hi i need help with these special factoring problems
5+^2-16=-12
x(x-11)=12
(x-6)(x+1)=-10
(2x+5)(x+1)
(a-4)(a+7)=-18

Answer by checkley77(12844)

5+^2-16=-12 believe you're missing a value preceeding the (^) sign?
---------------------
x(x-11)=12
x^2-11x-12=0
(x-12)(x+1)=0
x-12=0
1=12 ans
x+1=0
x=-1 ans
-------------------
(x-6)(x+1)=-10
x^2-5x-6+10=0
x^2-5x+4=0
(x-4)(x-1)=0
x-4=0
x=4 ans
x-1=0
x=1 ans.
-------------------------
(2x+5)(x+1) is there supposed to be an = sign here?
-------------------------------
(a-4)(a+7)=-18
a^2-4a+7a-28=-18
a^2+3a-28+18=0
a^2+3a-10=0
(a+5)(a-2)=0
a+5=0
a=-5 ans.
a-2=0
a=2 ans.